# Dynamics of “Spontaneous” Topic Changes in Next Token Prediction with Self-Attention

### v2

@@ Abstract
Human cognition can spontaneously shift conversation topics, often triggered by emotional or contextual signals. In contrast, self-attention-based language models depend on structured statistical cues from input tokens for next-token prediction, lacking this spontaneity. Motivated by this distinction, we investigate the factors that influence the next-token prediction to change the topic of the input sequence. We define concepts of topic continuity, ambiguous sequences, and change of topic, based on defining a topic as a set of token priority graphs (TPGs). Using a simplified single-layer self-attention architecture, we derive analytical characterizations of topic changes. Specifically, we demonstrate that (1) the model maintains the priority order of tokens related to the input topic, (2) a topic change can occur only if lower-priority tokens outnumber all higher-priority tokens of the input topic, and (3) unlike human cognition, longer context lengths and overlapping topics reduce the likelihood of spontaneous redirection. These insights highlight differences between human cognition and self-attention-based models in navigating topic changes and underscore the challenges in designing conversational AI capable of handling “spontaneous” conversations more naturally. To the best of our knowledge, no prior work has explored these questions with a focus as closely aligned to human conversation and thought.
@@ 1 Introduction
In human cognition, topic changes often occur sporadically, with little apparent structure or forethought: an spontaneous shift in focus during a conversation, a sudden leap between ideas when brainstorming, or an unexpected redirection in storytelling. These seemingly chaotic shifts are intrinsic to the human way of processing and communicating information (Christoff and Fox,, 2018; Mills et al.,, 2020). For instance, someone discussing their favorite book might suddenly mention a family gathering. This abrupt change may be due to an emotional connection, such as recalling reading that book during a family vacation, where sensory details like the scent of the ocean or the warmth of the sun trigger a vivid memory. The hippocampus helps link these sensory cues to memories, enabling a meaningful shift (Lisman et al.,, 2017). The prefrontal cortex then integrates these memories, emotions, and context to redirect conversations spontaneously (Miller and Cohen,, 2001).
However, large language models (LLMs) rely on contextual signals in the input to perform topic shifts. Unlike humans, who transition intuitively between topics, often driven by emotion or creativity, LLMs follow a structured, statistical approach, remaining on topic unless explicit cues signal a change. Figure 2 illustrates this distinction using the first sentence of the book “One hundred years of solitude” (García Márquez,, 1967). Additionally, related studies on conversational agents have recognized the practical challenges of spontaneous topic changes and proposed various practical approaches (Lim et al.,, 2010; Hwang et al.,, 2024; Xie et al.,, 2021; Soni et al.,, 2022; Ni et al.,, 2022; Lin et al.,, 2023).
It is also important to differentiate between topic changes and hallucinations. Hallucinations involve generating incorrect or fabricated information without clear contextual basis, often resulting in false or illogical content (Ji et al.,, 2023; Maynez et al.,, 2020). In contrast to topic changes, hallucinations are not deliberate or contextually driven shifts in discourse. More interestingly, some hallucinations may still appear to maintain topic consistency by staying loosely connected to the previous topic, yet they fail to introduce a meaningful or valid change. Our focus is on understanding when and why self-attention models predict next tokens from a different topic of the input, examining the conditions under which an input sequence from one topic leads to a prediction in another.
Recent advancements in the related field have substantially deepened our understanding of self-attention mechanisms in LLMs. The Transformer model, introduced by Vaswani et al., (2017), revolutionized natural language processing by using self-attention to determine the importance of each token in relation to others. This innovation enables Transformers to focus on relevant context when predicting the next token, significantly enhancing capabilities natural language understanding and generation(Brown et al.,, 2020; Clark et al.,, 2019; Devlin et al.,, 2019; Bisk et al.,, 2020). Additionally, connections between self-attention and support vector machines (SVMs) have offered valuable insights into optimization strategies, and mechanisms within the next token prediction (Tarzanagh et al., 2023b, ; Tarzanagh et al., 2023a, ; Li et al., 2024b, ). These advancements underscore the transformative role of attention-based architectures in token prediction tasks. A recent study by Li et al., (2023) highlights that in mixed topic inputs, transformers achieve higher pairwise attention between same-topic words compared to different-topic words. However, our approach aligns more closely with human cognitive processes, where the input is typically centered on a single topic, ensuring that discussions or conversations remain within a consistent context or theme.
@@
Figure 1: Illustration of the difference between human cognition and AI contextual completion. The original fragment of “One hundred years of solitude" (García Márquez,, 1967) (top) has a clear “spontaneous" thought, but the GPT-2’s completion [^2] Just to illustrate, we use the prompt Please continue this short sentence, forgetting about “One hundred Years of Solitude”, since on a real conversation the LLM would be blind to the final output. (bottom), demonstrates continuity.
Despite these advancements, our understanding of how LLMs manage topic changes compared to human cognition remains limited. Human cognition seamlessly navigates such transitions through spontaneous thought and contextual awareness (Christoff and Fox,, 2018). Investigating this distinction could provide valuable insights into the strengths and limitations of language models in mimicking human-like conversation. Bridging this gap is essential for enhancing language models, particularly in their ability to adapt to and manage dynamic topic changes–a process closely aligned with the flow of spontaneous thought in human cognition. Specifically, we address the following research questions:
  • What underlying mechanism drives topic changes during next token prediction in self-attention architectures?
  • Which factors shape the dynamics of topic changes in next token prediction?
To answer the above questions, it is essential first to explore how self-attention allocates attention to tokens for next token prediction with respect to the input topic. Understanding how attention weights are allocated offers key insights into the prioritization of tokens by self-attention mechanisms, particularly when models are trained on datasets spanning multiple topics. This fundamental understanding will pave the way for uncovering the role of attention patterns in driving “spontaneous” topic changes. To the best of our knowledge, there are no other studies that have investigated these dynamics so closely in relation to human conversation and thought processes.
@@
For simplicity, we elaborate our results using a two-topic scenario, but it is straightforward to extend them to multiple topics. Imagine an oracle that is an expert on Topic A, capable of following any conversation within that topic while staying true to its context. Now, suppose the oracle gains knowledge of Topic B and is following a conversation about Topic A. Will the oracle’s responses remain firmly within Topic A, or will the influence of the knowledge of Topic B cause the conversation to drift? This analogy encapsulates the problem we address: understanding when and why attention models might preserve a topic or change to another “spontaneously". Specifically, we make the following contributions:
@@
    Impact of sequence length and topic overlap. Theorem 5 establishes that longer input sequences decrease the likelihood of changing topics. Furthermore, overlapping topics act as a stabilizing factor, not increasing the frequency of “spontaneous” topic changes. This finding represents a key contribution to the development of human-like AI conversational models. Unlike human dialogue, where extended discussions often encourage spontaneous topic shifts and overlapping topics promote cognitive connections that facilitate smooth transitions, our results highlight a fundamental limitation of attention-based models in replicating these human-like dynamics.
Figure 2: Overview of our approach. We define a topic by a set of TPGs $\{\mathcal{G}^{(k)}\}^{K}_{k=1}$ (Def. 2) and generate a dataset for each topic such that the dataset can reconstruct the corresponding TPGs exactly (Assumption 4). The combination of dataset A and dataset B becomes the dataset for the mixed-topic model. We train self-attention models independently on each dataset. Then, we generate a new input sequence from topic A and predict the next token with two models, self-attention model A and self-attention model (mixed-topic). The next-token prediction with the mixed-topic model is categorized into three outcomes: keeps topic A (topic continuity from Def. 3), ambiguous sequence (from Def. 4), changes topic A (change of topic from Def. 5). Further details for each category are in Figure 4.
Following the framework of token-priority graphs (TPGs) introduced by Li et al., 2024b , we start with defining a topic as a set of TPGs (see Sec 3). For simplicity, we consider two different topics, Topic A and Topic B, characterized by two different sets of TPGs. We demonstrate the preservation of input topic priorities in a model trained with mixed-topic datasets (see Sec 4). Additionally, we introduce three concepts to categorize the prediction outputs: topic continuity, ambiguous sequence, and change of topic. Figure 2 outlines the structural overview of our approach. We subsequently establish the conditions under which a self-attention model induces a topic change, showing that the probability of topic changes does not increase with longer input sequences or the presence of overlapping topics (see Sec 5). All proofs are refer to Appendix A. To formalize this approach, we begin with the structured problem setup.
@@ 2 Problem Setup
Next Topic Prediction with Self-attention Model
@@
Token-priority Graph and Global Convergence of the Self-attention Model
Li et al., 2024b defined a token-priority graph (TPG) as a directed graph with nodes representing tokens in the vocabulary. $\text{DSET}^{(k)}$ is a subset of sequences from DSET with the same last token is ${\bf e}_{k}=\bar{\bf x}$. They defined TPGs $\{\mathcal{G}^{(k)}\}_{k=1}^{K}$ such that every $\mathcal{G}^{(k)}$ is a directed graph where for every sequence $({\bf X},y)\in\text{DSET}^{(k)}$ a directed edge is added from ${\bf e}_{y}$ to every token ${\bf x}\in{\bf X}$. TPGs are further divided into strongly-connected components (SCCs), which capture subsets of tokens where priority relations are cyclic, indicating comparable token importance. For tokens within two different SCCs, strict priority orders emerge, helping the model to differentiate between tokens when learning next-token predictions. We use the same notation as Li et al., 2024b , given a directed graph $\mathcal{G}$, for $i,j\in[K]$ such that $i\neq j$:
@@
  • $(i\Rightarrow j)\in\mathcal{G}$ denotes that the directed path $(i\rightarrow j)$ is present in $\mathcal{G}$ but $j\rightarrow i$ is not.
@@ 3 Defining Topics
In order to answer our research questions regarding the dynamics of topic changes we need to define the concept of a topic. In the previous settings, a dataset DSET generates TPGs $\{\mathcal{G}^{(k)}\}_{k=1}^{K}$, but, conversely, an existing set of TPGs can generate DSET, akin to generating data from a particular topic, which leads to the following definitions.
@@
We can define a topic as a set of TPGs because it inherently captures the relationships and contextual dependencies between tokens, offering a structured representation of a topic’s semantics. This means that we can use a topic $\mathbb{T}$ to generate DSET by sampling sequences $({\bf X},y)$ within $\mathbb{T}$. Given the finite number of edges, a DSET can be generated from $\mathbb{T}$ such that it can reconstruct the exact TPGs $\{\mathcal{G}^{(k)}\}_{k=1}^{K}$ that define $\mathbb{T}$, following the construction method in Li et al., 2024b . This leads to the following reasonable assumption:
@@
This definition establishes the criteria for a next predicted token to preserve the topic of the input sequence. Notice that if we have two topics, $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, we can generate two different datasets DSETa and DSETb. It is clear that $\tilde{\bf W}_{a}$ [^3] Notation: The subscripts of weights and objects correspond to the associated topic. For instance $\tilde{\bf W}_{a}$ denotes the weights defined in Equation 2, obtained from DSETa, which pertains to topic $\mathbb{T}_{a}$. trained only with DSETa will always keep topic $\mathbb{T}_{a}$. But we could also obtain $\tilde{\bf W}_{ab}$ with a dataset combining DSETa and DSETb as training sets. The central question is whether $\tilde{\bf W}_{ab}$ keeps topic $\mathbb{T}_{a}$, given an input sequence ${\bf X}$ that belongs to $\mathbb{T}_{a}$, or if it instead predicts tokens that prompt a topic change. This can be understood through an analogy: consider an oracle only proficient in Topic A, capable of following conversations in the context of Topic A. If this oracle gains knowledge of Topic B, will it continue to follow the conversation within Topic A’s context when the conversation is within Topic A, or will its response begin to exhibit a shift to Topic B?
@@ 4 Attention within Mixed Topics
Let’s first understand the way in which attention models assign priority to tokens within mixed-topic setting. For simplicity, we elaborate our results using a two-topic scenario, but it is straightforward to extend all of our theoretical results on multiple topics. Notice the self-attention embedding output is a linear combination of $\bf X$ given by $\mathbb{S}({\bf XW\bar{x}})$. The embeddings in $\bf X$ corresponding to the highest entries in the vector $\mathbb{S}({\bf XW\bar{x}})$ will receive higher priority in their importance to predict the next token, therefore we can hypothesize that models in which $\mathbb{S}({\bf XW\bar{x}})$ are ordered in a similar way will predict similar next tokens. This idea leads to our first main result which considers this situation within a mixed topics setting:
@@
[…]
Figure 3: Illustration of topic combination. $\mathcal{G}^{(4)}_{a}$ and $\mathcal{G}^{(4)}_{b}$ are token-priority graphs (TPGs) of the last token ${\bf e}_{4}$ for $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, respectively. $\mathcal{G}^{(4)}_{ab}$ combines the TPG for $\mathbb{T}_{a}$ and the TPG for $\mathbb{T}_{b}$.
@@
This implies that for an input sequence $\mathbf{X}$, a model trained in a mixed-topic setting will maintain the priority of the topic to which $\mathbf{X}$ belongs. Consequently, the attention will be allocated in the same order as if the model had been trained exclusively on the original topic of $\mathbf{X}$. The input sequence ${\bf X}=[{\bf e}_{5},{\bf e}_{1},{\bf e}_{3},{\bf e}_{4}]^{\top}$, as shown in Figure 4 (top), is generated from $\mathbb{T}_{a}$. The predicted next token $\hat{y}_{{\bf w}_{ab}}$ is ${\bf e}_{5}$ and the highest probability SCC in mixed-topics is $\widehat{\mathcal{G}}^{(4)}_{ab}({\bf X})=\{{\bf e}_{5}\}$. Since $\hat{y}_{{\bf w}_{ab}}$ belongs to ${\widehat{\mathcal{G}}}^{(4)}_{ab}({\bf X})$, ${\bf W}_{ab}$ for input sequence $\bf X$ is considered as topic continuity, based on the Definition 3.
@@
The only assumption about ${\bf X}$ on Theorem 3 is that it belongs to $\mathbb{T}_{a}$. However, if $\mathbf{X}$ belongs to $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, the priority will be preserved within both topics. Additionally, strict equality in the attention priority is maintained, but strict inequalities may not be when combining both datasets. In such cases, the union of their TPGs can form new SCCs, disrupting strict priority. As illustrated in Figure 3, we present $\mathcal{G}^{(4)}_{a}$ and $\mathcal{G}^{(4)}_{b}$ as the TPGs corresponding to last input token ${\bf e}_{4}$ for $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, respectively. In $\mathcal{G}^{(4)}_{a}$, the token priority is ${\bf e}_{5}>{\bf e}_{3}>{\bf e}_{1}={\bf e}_{2}>{\bf e}_{4}$. Once combining $\mathcal{G}^{(4)}_{a}$ and $\mathcal{G}^{(4)}_{b}$, the priority order in the mixed-topics is ${\bf e}_{5}>{\bf e}_{3}>{\bf e}_{1}={\bf e}_{2}={\bf e}_{4}$, based on $\mathcal{G}^{(4)}_{ab}$. The strict equality ${\bf e}_{1}={\bf e}_{2}$ from $\mathcal{G}^{(4)}_{a}$ is maintained in $\mathcal{G}^{(4)}_{ab}$, but the strict inequality ${\bf e}_{2}>{\bf e}_{4}$ is replaced with ${\bf e}_{2}={\bf e}_{4}$ in mixed-topics, forming the new SCC, $\{{\bf e}_{1},{\bf e}_{2},{\bf e}_{4}\}$, in $\mathcal{G}^{(4)}_{ab}$.
@@
Figure 4: Illustration of topic continuity, ambiguous sequence, and change of topic (see Definition 3,  4, and 5 for formal definitions). For each input sequence derived from $\mathbb{T}_{a}$, we employ a self-attention model trained on $\mathbb{T}_{a}$ and another model trained on the mixed-topics dataset $\mathbb{T}_{ab}$ to predict the next tokens $\hat{y}_{{\bf w}_{a}}$ and $\hat{y}_{{\bf w}_{ab}}$, respectively (orange). $\widehat{\mathcal{G}}^{(4)}_{ab}$ is the highest probability SCC (see Definition 1) in mixed-topics setting and $\widehat{\mathcal{G}}^{(4)}_{a}$ is the highest probability SCC in $\mathbb{T}_{a}$. Top: The input sequence is classified as topic continuity if $\hat{y}_{{\bf w}_{ab}}\in\widehat{\mathcal{G}}^{(4)}_{a}(X)$. Middle: The input sequence is ambiguous if $\hat{y}_{{\bf w}_{ab}}\not\in\widehat{\mathcal{G}}^{(4)}_{a}(X)$, but $\widehat{\mathcal{G}}^{(4)}_{a}(X)\subset\widehat{\mathcal{G}}^{(4)}_{ab}(X)$. Bottom: ${\bf W}_{ab}$ is considered to change of topic, if $\hat{y}_{{\bf w}_{ab}}\not\in\widehat{\mathcal{G}}^{(4)}_{a}(X)$ and $\widehat{\mathcal{G}}^{(4)}_{a}(X)\not\subset\widehat{\mathcal{G}}^{(4)}_{ab}(X)$. The numeric details are provided in Appendix C.3.
The attention priority similarity quantifies how well the weights ${\bf W}^{\prime}$ preserve the attention priority of the weights ${\bf W}$. A value of 1 indicates that the priority is fully preserved. Using this metric, we conduct experiments, with results in Figure 5. We generate embeddings with $K=10$ and $d=16$, and randomly construct TPGs for $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$. Using these TPGs, we randomly generate $\text{DSET}_{a}$ and $\text{DSET}_{b}$. We compute $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$, $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$ and $\frac{{\bf W}_{ab}(\tau)}{\|{\bf W}_{ab}(\tau)\|_{F}}$ using the same procedure as Li et al., 2024b . We generate test sequences ${\bf Z}$ within $\mathbb{T}_{a}$, and we calculate the attention priority similarity of $\frac{{\bf W}_{ab}(\tau)}{\|{\bf W}_{ab}(\tau)\|_{F}}$ relative to both $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ and $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$. We repeat this process for multiple TPGs and input sequences (simulation details in Appendix C). Figure 5 clearly demonstrates that the similarity converges to 1 after $\tau=8000$ iterations when evaluated relative to $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ (blue line), but fails to converge relative to $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$ (orange line). These observations align with the results of Theorem 3.
Figure 5: Convergence of attention priority similarity for $\frac{{\bf W}_{ab}(\tau)}{\|{\bf W}_{ab}(\tau)\|_{F}}$ relative to $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ (blue) and $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ (orange).

### v3

# Dynamics of Spontaneous Topic Changes in Next Token Prediction with Self-Attention

@@ Abstract
Human cognition is punctuated by abrupt, spontaneous shifts between topics—driven by emotional, contextual, or associative cues—a phenomenon known as spontaneous thought in neuroscience. In contrast, self-attention based models depend on structured patterns over their inputs to predict each next token, lacking spontaneity. Motivated by this distinction, we characterize spontaneous topic changes in self-attention architectures, revealing both their similarities and their divergences from spontaneous human thought. First, we establish theoretical results under a simplified, single-layer self-attention model with suitable conditions by defining the topic as a set of Token Priority Graphs (TPGs). Specifically, we demonstrate that (1) the model maintains the priority order of tokens related to the input topic, (2) a spontaneous topic change can occur only if lower-priority tokens outnumber all higher-priority tokens of the input topic, and (3) unlike human cognition, the longer context length or the more ambiguous input topic reduces the likelihood of spontaneous change. Second, we empirically validate that these dynamics persist in modern, state-of-the-art LLMs, underscoring a fundamental disparity between human cognition and AI behaviour in the context of spontaneous topic changes. To the best of our knowledge, no prior work has explored these questions with a focus as closely aligned to human thought.
@@ 1 Introduction
Human cognition is punctuated by abrupt, apparently unstructured topic changes, the hallmark of human spontaneous thought, a phenomenon that has become a central topic in cognitive neuroscience (Bellana et al.,, 2022; Christoff and Fox,, 2018; Christoff et al.,, 2011; Kucyi et al.,, 2023; Mildner and Tamir,, 2019, 2024; Mills et al.,, 2020). For example, a spontaneous shift in focus during a conversation, a sudden leap between ideas when brainstorming, or an unexpected redirection in storytelling. These abrupt changes may be due to an emotional connection, such as recalling reading that book during a family vacation, where sensory details like the scent of the ocean or the warmth of the sun trigger a vivid memory. However, LLMs rely on contextual signals in the input to perform topic shifts, instead of spontaneous topic changes. They follow a structured, statistical approach, remaining on topic unless explicit cues signal a change. Figure 2 illustrates this distinction using the first sentence of the book “One hundred years of solitude” (García Márquez,, 1967).
Our work takes initial steps toward formalizing the dynamics of spontaneous topic changes in LLMs and analyzing how they relate to or diverge from human spontaneous thought. To this end, we ground our theoretical analysis in a single-layer self-attention model and empirically extend it to modern LLMs, laying groundwork for drawing parallels between AI models and human cognition.
Recent advancements in the related field have substantially deepened our understanding of self-attention architectures. (Li et al., 2024b, ; Tarzanagh et al., 2023a, ; Tarzanagh et al., 2023b, ) have linked the self-attention to support vector machines (SVMs), offering optimization strategies for next-token prediction. A study by Li et al., (2023) highlights that in mixed-topics inputs, transformers achieve higher pairwise attention between same-topic words compared to different-topic words. In parallel, prior studies have recognized the practical challenges of spontaneous topic changes in LLMs and proposed various approaches to address them (Hwang et al.,, 2024; Lim et al.,, 2010; Lin et al.,, 2023; Ni et al.,, 2022; Soni et al.,, 2022; Xie et al.,, 2021). Notably, spontaneous topic changes must be differentiated from hallucinations, generating incorrect or fabricated information without clear contextual basis (Ji et al.,, 2023; Maynez et al.,, 2020).
Despite these advancements, our understanding of the dynamics of spontaneous topic changes in LLMs remains limited. Investigating the parallels and differences between spontaneous topic changes in self-attention models and human spontaneous thought could provide valuable insights for improving the efficiency of language models with human-like thought. Since modern LLMs rely on self-attention architectures, we begin by theoretically characterizing spontaneous topic shifts in a simplified setting. We then extend these findings through experiments on more complex, state-of-the-art models. To the best of our knowledge, no prior studies have investigated these dynamics so closely in relation to human thought.
Figure 2 outlines our theoretical framework. To make the mathematical analysis tractable, we follow the same single-layer self-attention framework with log-loss objective function governed by Assumptions 14 from Li et al., 2024b . Inspired by token-priority graphs (TPGs) (Li et al., 2024b, ) and building on attribution graphs from Ameisen et al., (2025) for exposing an LLM’s internal computation, we define a topic as a set of TPGs. This graph-based formulation aligns naturally with recent advances in structured representations for LLMs (Sen et al.,, 2023; Wang et al.,, 2025). Furthermore, this mirrors neuroscience models of human spontaneous thought, in which concepts serve as nodes connected by associative edges (Mildner and Tamir,, 2019). Despite relying on these specific settings, our experiments extend our findings to modern LLMs, empirically confirming that relaxing these assumptions does not seem to undermine our core insights.
@@
Figure 1: Illustration of the difference between human cognition and LLMs. The original fragment of “One hundred years of solitude" (García Márquez,, 1967) (top) has a clear spontaneous thought, but the GPT-2’s completion (bottom), demonstrates continuity. [^2] Just to illustrate, we use the prompt Please continue this short sentence, forgetting about “One hundred Years of Solitude”, since on a real conversation the LLM would be blind to the final output.
Figure 2: Overview of our theoretical framework. We define a topic by a set of TPGs $\{\mathcal{G}^{(k)}\}^{K}_{k=1}$ (Def. 2) and generate a dataset for each topic. The combination of dataset A and dataset B becomes the dataset for the mixed-topic model. We train self-attention models independently on each dataset. Then, we generate a new input sequence from topic A and predict the next token with two models, self-attention model A and self-attention model (mixed-topic). The next-token prediction with the mixed-topic model is categorized into three outcomes: keeps topic A (topic continuity from Def. 3), ambiguous sequence (from Def. 4), changes topic A (change of topic from Def. 5). Further details for each category are in Figure 3.
@@
Imagine an oracle that is an expert on Topic A, capable of following any conversation within that topic while staying true to its context. Now, suppose the oracle gains knowledge of Topic B and is following a conversation about Topic A. Will the oracle’s responses remain within Topic A, or will the influence of the knowledge of Topic B cause the conversation to drift? This analogy encapsulates the problem we address: understanding when and why attention models might preserve a topic or change to another spontaneously. Specifically, we make the following contributions:
@@
    Impact of sequence length and topic ambiguity. Theorem 5 establishes that longer input sequences decrease the likelihood of changing topics. Furthermore, topic ambiguity acts as a stabilizing factor, not increasing the frequency of spontaneous topic changes.
  4. 4.
    Difference between LLMs and human cognition. In Section 6 we empirically extend Theorem 5 to modern deeper LLMs. Unlike human cognition, where extended discussions often encourage spontaneous thoughts and topic ambiguity promotes cognitive connections, our results highlight an opposite behavior in LLMs: neither longer prompts nor greater topic ambiguity appreciably raise the likelihood of an spontaneous topic change.
Overview of the paper structure. We begin with the problem setup in Sec 2. Sec 3 introduces the definition of topic, and Sec 4 examines how self-attention models allocate the token priorities within the mixed-topics. In Sec 5, we establish the conditions under which a self-attention model induces spontaneous topic changes and show the dynamics of topic changes with longer input sequences or the presence of topic ambiguity. We then extend our analysis to frontier LLMs in Sec 6. Related work and discussion are provided in Secs 7 and 8, respectively. All proofs are provided in Appendix A.
@@ 2 Problem Setup
2.1 Next Topic Prediction with Self-attention Model
@@
2.2 Token-priority Graph and Global Convergence of the Self-attention Model
Li et al., 2024b defined a token-priority graph (TPG) as a directed graph with nodes representing tokens in the vocabulary. $\text{DSET}^{(k)}$ is a subset of sequences from DSET with the same last token is ${\bf e}_{k}=\bar{\bf x}$. They defined TPGs $\{\mathcal{G}^{(k)}\}_{k=1}^{K}$ such that every $\mathcal{G}^{(k)}$ is a directed graph where for every sequence $({\bf X},y)\in\text{DSET}^{(k)}$ a directed edge is added from ${\bf e}_{y}$ to every token ${\bf x}\in{\bf X}$. TPGs are further divided into strongly-connected components (SCCs), which capture subsets of tokens with equal priority. For tokens within two different SCCs, strict priority orders emerge, helping the model to differentiate between tokens when learning next-token predictions. We use the same notation as Li et al., 2024b , given a directed graph $\mathcal{G}$, for $i,j\in[K]$ such that $i\neq j$:
@@ 2.2 Token-priority Graph and Global Convergence of the Self-attention Model
  • $(i\Rightarrow j)\in\mathcal{G}$ denotes that the directed path $(i\rightarrow j)$ is presented in $\mathcal{G}$ but $j\rightarrow i$ is not.
@@ 3 Defining Topics
In order to answer our research questions regarding the dynamics of topic changes we need to define the concept of a topic. In the previous settings, a dataset DSET generates TPGs $\{\mathcal{G}^{(k)}\}_{k=1}^{K}$, but, conversely, an existing set of TPGs can generate DSET. Therefore, inspired by Ameisen et al., (2025) that introduces attribution graphs to reveal the LLMs’ internal computational structure, we define a topic as a set of TPGs:
@@
Our graph-based formulation aligns with recent advances in structured representations of LLMs (Sen et al.,, 2023; Wang et al.,, 2025). Given the finite number of edges, a DSET can be generated from $\mathbb{T}$ such that it can reconstruct the exact TPGs $\{\mathcal{G}^{(k)}\}_{k=1}^{K}$ that define $\mathbb{T}$, following the construction method in Li et al., 2024b . This leads to the following reasonable assumption:
@@
Remark. If we have two topics, $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, we can generate two different datasets DSETa and DSETb. The union of $\{\mathcal{G}^{(k)}_{a}\}_{k=1}^{K}$ for $\mathbb{T}_{a}$ and $\{\mathcal{G}^{(k)}_{b}\}_{k=1}^{K}$ for $\mathbb{T}_{b}$ forms $\{\mathcal{G}^{(k)}_{ab}\}_{k=1}^{K}$ for the mixed-topics, denoted as the TPGs for the combined DSETa and DSETb.
It is clear that $\tilde{\bf W}_{a}$ trained only with DSETa will always keep topic $\mathbb{T}_{a}$. [^3] Notation: The subscripts of weights and objects correspond to the associated topic. For instance $\tilde{\bf W}_{a}$ denotes the weights defined in Equation 2, obtained from DSETa, which pertains to topic $\mathbb{T}_{a}$. But we could also obtain $\tilde{\bf W}_{ab}$ with a dataset combining DSETa and DSETb as training sets. The central question is whether $\tilde{\bf W}_{ab}$ keeps topic $\mathbb{T}_{a}$, given an input sequence ${\bf X}$ that belongs to $\mathbb{T}_{a}$, or if it instead predicts tokens that prompt a topic change.
@@ 4 Attention within Mixed Topics
Let’s first understand the way in which attention models assign priority to tokens within mixed-topic setting. For simplicity, we elaborate our results using a two-topic scenario, but it is straightforward to extend the results on multiple topics. Notice the self-attention embedding output is a linear combination of $\bf X$ given by $\mathbb{S}({\bf XW\bar{x}})$. The embeddings in $\bf X$ corresponding to the highest entries in $\mathbb{S}({\bf XW\bar{x}})$ will receive higher priority to predict the next token, therefore we can hypothesize that models in which $\mathbb{S}({\bf XW\bar{x}})$ are ordered in a similar way will predict similar next tokens. This idea leads to our first main result which considers this situation within a mixed topics setting:
@@
  • $\text{if }\quad[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{i}=[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{j}\text{, then}\quad[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{i}=[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{j}$
  • $\text{if }\quad[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{i}>[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{j}\text{, then}\quad[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{i}\geq[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{j}$
  • $\text{if }\quad[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{i}<[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{j}\text{, then}\quad[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{i}\leq[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{j}$
@@
This implies that for an input sequence $\mathbf{X}$, a model trained in a mixed-topic setting will maintain the priority of the topic to which $\mathbf{X}$ belongs. Consequently, the attention will be allocated in the same order as if the model had been trained exclusively on the original topic of $\mathbf{X}$. For the first input sequence ${\bf X}=[{\bf e}_{5},{\bf e}_{1},{\bf e}_{3},{\bf e}_{4}]^{\top}$ from $\mathbb{T}_{a}$, as shown in Figure 3 (right), the predicted next token $\hat{y}_{{\bf w}_{ab}}$ is ${\bf e}_{5}$ and the highest probability SCC in mixed-topics is $\widehat{\mathcal{G}}^{(4)}_{ab}({\bf X})=\{{\bf e}_{5}\}$. Since $\hat{y}_{{\bf w}_{ab}}$ belongs to ${\widehat{\mathcal{G}}}^{(4)}_{ab}({\bf X})$, ${\bf W}_{ab}$ for input sequence $\bf X$ is considered as topic continuity, based on the Definition 3.
@@
The only assumption about ${\bf X}$ on Theorem 3 is that it belongs to $\mathbb{T}_{a}$. However, if $\mathbf{X}$ belongs to $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, the priority will be preserved within both topics. Additionally, strict equality in the attention priority holds, but strict inequalities may not, as the union of their TPGs can form new SCCs, like $\{{\bf e}_{1},{\bf e}_{2},{\bf e}_{4}\}$ in $\mathcal{G}^{(4)}_{ab}$ (left side of Figure 3), disrupting strict priority (see details in Appendix C).
@@
Figure 3: Depiction of each scenario in next token prediction. Left: Taking the last token ${\bf e}_{4}$ as an example, $\mathcal{G}^{(4)}_{ab}$ for $\mathbb{T}_{ab}$ is formed by the union of $\mathcal{G}^{(4)}_{a}$ and $\mathcal{G}^{(4)}_{b}$. Right: For each input sequence belonging to $\mathbb{T}_{a}$, we employ a self-attention model trained on $\mathbb{T}_{a}$ and another model trained on the mixed-topics dataset $\mathbb{T}_{ab}$ to predict the next tokens, denoted as $\hat{y}_{{\bf w}_{a}}$ and $\hat{y}_{{\bf w}_{ab}}$ (orange). $\widehat{\mathcal{G}}^{(4)}_{ab}$ and $\widehat{\mathcal{G}}^{(4)}_{a}$ represent the highest probability SCCs (Definition 1) in mixed-topics setting and in $\mathbb{T}_{a}$, respectively. There are three scenarios, topic continuity (Definition 3), ambiguous sequence (Definition 4), and change of topic (Definition 5). The numeric details for each scenario are provided in Appendix D.5.